OpenStudy (anonymous):

If 0 lies in the quadrant IV. What can be the value of cos 0 ?

OpenStudy (january123):

What do you think?

OpenStudy (anonymous):

The value of all cos values is positive in the fourth quadrant right?

OpenStudy (solomonzelman):

cos(0)=cos(45-45)=sin(45)cos(45)+sin(45)cos(45)